OpenStudy (cypress7):

can someone help me please? which fractions is equivalent to -3/-8 Choose all answers that are correct. A. -6/-16 B.12/32 C.-3/8 D.9/-24

OpenStudy (cypress7):

wait one second I messed up it should be what's equivalent to -3/-8

OpenStudy (danjs):

ok

OpenStudy (danjs):

You are just multiplying the fraction by 1. -3/-3 = 1

OpenStudy (danjs):

oh you changed the question

OpenStudy (cypress7):

okay so 3-/8 is correct right? and I know -6/-16 is right too

OpenStudy (danjs):

\[\frac{ -3 }{ -8 }\frac{ -4 }{ -4 } = \frac{ 12 }{ 32 }\]

OpenStudy (danjs):

Multiplying top and bottom of -3/-8 by -4 gives 12/32

OpenStudy (cypress7):

okay so A B and C?

OpenStudy (danjs):

no not c... -3/-8 = +3/8

OpenStudy (danjs):

negative divided by negative is a positive number

OpenStudy (cypress7):

oh sorry Im not good at fractions

OpenStudy (danjs):

Just A and B A is a fraction of 2/2 times the original B is a fraction of (-4)/(-4) time the original.

OpenStudy (cypress7):

thank you danJS! you helped a lot!

OpenStudy (danjs):

np